Please help me solve this equation: X^5/X^2
Answer by anantha(86)

sol:
in x^5 the base=x and exponent is 5
in x^2 the base =x and exponent is 2
here bases are same so we can subtract the exponents
x^5/x^2= x^(5-2)=x^3

2nd method
in x^5 the base=x and exponent is 5 , x^5 means we can multiply the base 5 times
means
x^5=x*x*x*x*x
in x^2 the base=x and exponent is 2,x^2 means we can multiply the base 2 times
means
x^2=x*x
now
x^5/x^2=(x*x*x*x*x)/(x*x)
after cancellation we have x*x*x
x^5/x^2=x*x*x=x^3